OpenStudy (anonymous):

13x-(2y+1)+12y-2+4(x+2y-16)-22=3(2x+12y)-9x+2-6y-2(y+3)

OpenStudy (anonymous):

\[-25+13 x+10 y+4 (-16+x+2 y)= 2-9 x-6 y-2 (3+y)+3 (2 x+12 y)\]Expand and simplify the LHS and the RHS.\[17 x+18 y-89=-3 x+28 y-4 \]Move the RHS to the LHS and simplify.\[20 x-10 y-85=0\]